OpenStudy (anonymous):

this is hard '.' Which improper fraction is equivalent to two and three over nine? fourteen over nine twenty-nine over nine twenty-one over nine fifteen over nine

OpenStudy (zzr0ck3r):

2 and 3 over 9?

OpenStudy (anonymous):

2 3/9

OpenStudy (zzr0ck3r):

ok lets look at the first one \[\frac{14}{9}\] my mom just told me we are leaving so I will do this one and you can figure it out how many times does 9 go into 14 once what is left over? 5 so \[\frac{14}{9}=1\frac{5}{9}\]

OpenStudy (zzr0ck3r):

do the same thing for the rest to find the answer

OpenStudy (texaschic101):

2 3/9 -- turn to improper fraction by taking the whole number (2) and multiplying it by the denominator (9), then adding that to the numerator (3), then putting it all over the original denominator (9). So, 2 3/9 = (2 x 9 + 3) / 9....you can do the math